### v2

# Ram-EHR: Retrieval Augmentation Meets Clinical Predictions on Electronic Health Records

Abstract
We present Ram-EHR, a Retrieval AugMentation pipeline to improve clinical predictions on Electronic Health Records (EHRs). Ram-EHR first collects multiple knowledge sources, converts them into text format, and uses dense retrieval to obtain information related to medical concepts. This strategy addresses the difficulties associated with complex names for the concepts. Ram-EHR then augments the local EHR predictive model co-trained with consistency regularization to capture complementary information from patient visits and summarized knowledge. Experiments on two EHR datasets show the efficacy of Ram-EHR over previous knowledge-enhanced baselines (3.4% gain in AUROC and 7.2% gain in AUPR), emphasizing the effectiveness of the summarized knowledge from Ram-EHR for clinical prediction tasks. The code will be published at https://github.com/ritaranx/RAM-EHR.
Ram-EHR: Retrieval Augmentation Meets Clinical Predictions
on Electronic Health Records
Ran Xu1thanks:  Equal contribution., Wenqi Shi211footnotemark: 1, Yue Yu2, Yuchen Zhuang2, Bowen Jin3 May D. Wang2, Joyce C. Ho1, Carl Yang1 1 Emory University  2 Georgia Institute of Technology 3 University of Illinois at Urbana Champaign {ran.xu,joyce.c.ho,j.carlyang}@emory.edu, {wshi83,yczhuang,yueyu,maywang}@gatech.edu, bowenj4@illinois.edu
1 Introduction
Electronic Health Records (EHRs), encompassing detailed information about patients such as symptoms, diagnosis, and medication, are widely used by physicians to deliver patient care (Jensen et al., 2012; Cowie et al., 2017; Shi et al., 2024). Recently, a vast amount of deep learning models have been developed on EHR data (Choi et al., 2016, 2020; Gao et al., 2020; Xu et al., 2022; Cai et al., 2022; Wang et al., 2023b) for various downstream prediction tasks (e.g., disease diagnosis, risk prediction) to facilitate precision healthcare.
To further improve the downstream predictive performance, several works attempt to augment the EHR visits with external knowledge. For example, van Aken et al. (2021) and Naik et al. (2022) incorporate additional clinical notes, although these clinical notes can be noisy and contain irrelevant contents for clinical predictions; another solution is to leverage external clinical knowledge graphs (KGs), such as UMLS (Chandak et al., 2023), which contains rich medical concepts (e.g., disease, medications) and their corresponding relationships (Jin et al., 2023). Integrating KGs with EHRs has been shown to boost model performance (Choi et al., 2017; Xu et al., 2023b; Gao et al., 2023). However, these works mostly rely on knowledge from a single source and medical KGs mainly focus on specific types of relations (e.g., hierarchical relations), which do not comprehensively capture the semantic information for medical codes (e.g., phenotype). Besides, it is non-trivial to align medical codes in EHRs with KGs due to the non-uniformity of surface names (e.g., abbreviations or colloquial terms) (Hao et al., 2021; Zhang et al., 2022). Jiang et al. (2024) use knowledge generated from large language models (LLMs) to assist EHR prediction, but LLMs may not always provide the most relevant knowledge for target tasks and face the risk of hallucination. Effectively leveraging external knowledge to facilitate EHR predictive tasks remains a significant challenge.
In this work, we propose Ram-EHR, a retrieval-augmented framework tailored for clinical predictive tasks on EHRs. Instead of leveraging a single knowledge source, Ram-EHR collects multiple knowledge sources (e.g., KGs, scientific literature) and converts them to text corpus, which enjoys the merits of a more comprehensive coverage of knowledge in a unified format. Then, to obtain unified representations for different knowledge sources, we leverage dense retrieval (DR) (Karpukhin et al., 2020; Izacard et al., 2022; Yu et al., 2022b; Lin et al., 2023) to encode corpus and medical codes as dense vectors, intuitively capturing the semantics of medical codes and addressing the alignment issue between EHR and external knowledge. Finally, to reduce irrelevant information and reduce the length of the input text, we utilize an LLM to summarize the top-retrieved passages into concise and informative knowledge summaries relevant to downstream tasks for each medical code. This process enhances the relevance and utility of the retrieved knowledge for clinical tasks.
To leverage external knowledge to assist clinical prediction, we introduce a retrieval-augmented model alongside the local EHR predictive model, which relies solely on patient visit information. The augmented model concatenates summarized passages and medical codes, feeding them into a moderate-size, pre-trained language model. We then co-train the local model and the augmented model with a consistency regularization, which captures the complementary information from patient visits and summarized knowledge and helps the model with better generalization (Wan, 2009; Lang et al., 2022).
We verify the effectiveness of Ram-EHR by conducting experiments on two EHR datasets and show that Ram-EHR outperforms strong knowledge-enhanced predictive baselines by 3.4% in AUROC and 7.2% in AUPR on average. Our analysis further confirms the advantage of leveraging multi-source external knowledge as well as retrieval augmentation as plugins to assist vanilla EHR predictive models based on visits only.
Our contribution can be summarized as
  • We introduce Ram-EHR, an innovative retrieval-augmented framework designed to harness external knowledge to enhance EHR-based clinical predictions. Notably, Ram-EHR offers flexibility and can seamlessly integrate diverse sources of knowledge.
  • We introduce a co-training approach aimed at effectively leveraging the complementary information derived from summarized knowledge and patient visit records.
  • We conduct experiments on two EHR datasets and observe that Ram-EHR leads to consistent gains when compared with other knowledge-enhanced EHR predictive models. We also demonstrate that Ram-EHR can serve as a generic plugin for multiple types of base models. Additional human studies justify the usefulness of summarized knowledge for assisting clinical prediction tasks.
2 Related Works
Retrieval Augmented Learning and its Application in Clinical Domain.
Retrieval augmented learning, which collects additional contextual information from external corpus, has shown effectiveness on diverse tasks including language modeling (Borgeaud et al., 2022; Wang et al., 2023a), knowledge-intensive NLP (Lewis et al., 2020; Shi et al., 2023a), commonsense reasoning (Yu et al., 2022a), code genereation (Parvez et al., 2021), and few/zero-shot learning (Xu et al., 2023a; Izacard et al., 2023). Compared to the general domain, the application of retrieval-augmented learning to clinical tasks is still under-explored. Some efforts have been paid to retrieval augmented clinical language models (Zakka et al., 2024), clinical decision making (Shi et al., 2023b; Thompson et al., 2023) and clinical question answering (Wang et al., 2023c; Xiong et al., 2024; Jeong et al., 2024). The most relevant works are Ye et al. (2021); Naik et al. (2022), which leverage clinical literature to augment clinical predictive models. Compared to these works, our contribution lies in two folds: (1) we design a retrieval augmentation for structured EHRs with a diverse collection of external knowledge, which provides more relevant information for target clinical prediction tasks; (2) we incorporate a co-training scheme to leverage both the visit-level information and external knowledge for predictions.
Knowledge-enhanced EHR Predictive Models.
Many studies attempt to harness external knowledge for clinical prediction tasks. The majority of them leverage structured knowledge, such as medical ontology (Choi et al., 2017; Panigutti et al., 2020; Lu et al., 2021), to capture hierarchical relationships among medical codes, or employ personalized knowledge graphs (Xu et al., 2023b; Jiang et al., 2024) to integrate patient-specific information. However, these methods often suffer from limited coverage of all medical codes due to the complexity of surface names. Alternatively, some approaches utilize unstructured medical text for health prediction tasks (Ye et al., 2021). However, Ye et al. (2021) rely on a restricted corpus of approximately 30,000 passages as their external corpus, resulting in limited coverage. Besides, it only considers ICD codes and uses the Levenshtein distance for retrieval, which cannot fully capture the semantics of medical codes.
3 Methodology
Figure 1: An overview of retrieval augmentation framework (left) and a detailed workflow of Ram-EHR (right). Ram-EHR initially gathers multiple knowledge sources and converts them into textual format. We then use dense retrieval to obtain information related to medical concepts. Next, we design an additional module to augment the local EHR predictive model co-trained with consistency regularization, capturing complementary information from both patient visits and summarized knowledge.
Figure 1 presents a workflow of Ram-EHR, with a specific focus on dense retrieval from multiple knowledge sources and consistency regularization with co-training schema. The details of Ram-EHR will be introduced in the following sections.
3.1 Problem Setup
The EHR data consists of a group of patients $\mathcal{P}$ with corresponding hospital visits $V=\{v_{1},v_{2},...,v_{|V|}\}$. Each visit $v_{i}$ includes a set of medical codes $C_{i}\subset\mathcal{C}$, where $\mathcal{C}$ is the total set of medical codes for $\mathcal{P}$. In this study, $\mathcal{C}$ contains multiple types of medical codes including diseases, medications, and procedures. Each medical code $c_{i}\in C_{i}$ is a clinical concept, and it is associated with a name $s_{i}$ in the form of short text snippets. Given the clinical record $v_{i}$ with the involved medical codes $\mathcal{C}_{i}$, we aim to predict the patient’s clinical outcome $y_{i}$ (a binary label).
3.2 Retrieval Augmentation w/ Medical Codes
Existing approaches often treat each visit as context-free vectors, which fail to capture the concrete semantics of medical codes. Being aware of this, we aim to create the summarized knowledge for each medical code $c_{i}$ using its surface name $s_{i}$ via retrieval augmentation with additional contexts.
Multi-source Corpus Creation. Retrieval augmentation requires additional corpora as external knowledge. To ensure the coverage of clinical knowledge, we collect a diverse external resources =$\mathcal{M}=\{d_{1},d_{2},\ldots,d_{|\mathcal{M}|}\}$. We represent each knowledge unit as a raw text to facilitate retrieval. The detailed information of $\mathcal{M}$ is in Appendix B.
Passage Retrieval. Given a collection of $|\mathcal{M}|$ passages, the objective of the retriever is to transform passages in a dense vector, so that it can efficiently retrieve the most relevant information to the input query. In our work, we adopt Dragon (Lin et al., 2023), a dual-encoder model with strong performance across domains as the retriever. Specifically, we first use the passage encoder $R_{D}(\cdot)$ to build an index for corpus $\mathcal{M}$ to support retrieval. Then, at runtime, we use the query encoder $R_{Q}(\cdot)$ to map the input to an embedding (same dimension as the passage embedding) and calculate the similarity as $f(q,d)={R}_{Q}(q)^{\top}{R}_{D}(d)$. For the medical code $c_{i}$ with the surface name $s_{i}$, we retrieve top-$k$ ($k=5$ in this work) passages ${\mathcal{T}}_{i}$ from the corpus $\mathcal{M}$ as
${\mathcal{T}}_{i}=\underset{d\in\mathcal{M}}{\operatorname{Top-}k}\ f(s_{i},d).$ (1)
The top retrieved passages are considered as the external knowledge for the medical code $c_{i}$.
Summarized Knowledge Generation. Although ${\mathcal{T}}_{i}$ contains the most relevant information for $c_{i}$ from $\mathcal{M}$, directly using them to assist predictions can be suboptimal, as simply concatenating these passages often leads to long contexts, and some of the retrieved passages can also be irrelevant (Yu et al., 2023). Motivated by the fact that LLMs have strong capabilities in text summarization (Zhang et al., 2024), we propose to use the off-the-shelf LLM (gpt-3.5-turbo-0613) to generate the summarized knowledge $e_{i}$ for medical code $c_{i}$ as
$e_{i}=\operatorname{LLM}([\text{Prompt},t_{i,1},\cdots,t_{i,k}]),$ (2)
where $t_{i}\in{\mathcal{T}}_{i}$ stands for the retrieved passages in Eq.(1). We incorporate information related to the downstream task within our prompt to ensure the generated summaries are task-specific. Detailed prompt designs can be found in Appendix D.
Remark. The retrieval step is efficient as the corpus indexing only needs to be done once before applying to prediction tasks. It only needs one extra ANN retrieval operation per query, which is efficiently supported by FAISS Johnson et al. (2021). Besides, we cache the summarized knowledge for each medical code to avoid redundant operations.
Algorithm 1 Overview of Ram-EHR.
1:  Input: $\mathcal{P}$: patients; $V$: corresponding hospital visits of patients.
2:  Initializing multi-source external knowledge $\mathcal{M}$;
3:  for $i=1,\cdots,|V|$ do
4:     for $c_{i}\in v_{i}$ do
5:        Get the medical code $c_{i}$ and the corresponding textual name $s_{i}$ included in visit $v_{i}$;
6:        // Passage Retrieval
7:        Retrieve passages ${\mathcal{T}}_{i}$ via Eq. (1)
8:        // Knowledge Summarization (Accelerated with caching)
9:        Summarize knowledge $e_{i}$ for $c_{i}$ via Eq. (2);
10:     end for
11:     // Co-training
12:     Predict $\widehat{y}_{i,1}$ with knowledge-augmented model $g_{\phi}$ via Eq. (3);
13:     Predict $\widehat{y}_{i,2}$ with visit-based local model $f_{\theta}$ via Eq. (4);
14:     // Update Model Parameters
15:     Compute loss function $\mathcal{L}$ via Eq. (5);
16:     Update model parameters $\phi$ and $\theta$;
17:  end forOutput: Augmented model $g_{\phi}$ and local model $f_{\theta}$; Final prediction $\widetilde{y}_{j}=\beta\widehat{y}_{j,1}+(1-\beta)\widehat{y}_{j,2}$ for the $j$-th test example $p_{j}$.
3.3 Augmenting Patient Visits with Summarized Knowledge via Co-training
Recall that patient visits and summarized knowledge encode complementary information for clinical prediction tasks — visits capture co-occurrence relationships, while summarized knowledge encodes semantic information. To effectively aggregate these two types of information, we design a co-training approach, detailed as follows.
Augmented Model $g_{\phi}$ with Summarized Knowledge. For patient $p_{i}$ having the hospital visit $v_{i}$ with involved medical codes $\mathcal{C}_{i}$, we decompose $\mathcal{C}_{i}$ into three subsets: $\mathcal{C}_{i}^{\operatorname{d}}$ for diseases, $\mathcal{C}_{i}^{\operatorname{m}}$ for medications, and $\mathcal{C}_{i}^{\operatorname{p}}$ for procedures. For each type of medical code, we flatten the visit into a document by concatenating all the codes and their summarized knowledge in a reversed sequential order. For example, for disease code $\mathcal{C}_{i}^{\operatorname{d}}$, the flattened document can be $X^{\operatorname{d}}_{i}=\left\{[\operatorname{CLS}],D_{t},D_{t-1},\ldots,D_{1}\right\}$, where $D_{i}=||_{c\in D_{i}}(c,e)$ is the concatenation of disease code and its summarized knowledge (Eq. 2) within the $i$-th visit. We then use a pre-trained language model (PLM) with a multi-layer perceptron (MLP) classification head as $g_{\phi}$ for prediction with flattened documents as inputs:
$\bm{h}_{i}^{k}=\operatorname{PLM}(X_{i}^{k}),~{}~{}\widehat{y}_{i,1}=\operatorname{MLP}\left(||_{k\in{\mathcal{S}}}\bm{h}_{{i}}^{{k}}\right).$ (3)
Here ${\mathcal{S}}=\{\operatorname{p},\operatorname{m},\operatorname{d}\}$, $\bm{h}_{i}$ is the representation of [CLS] token of $X_{i}$, $\widehat{y}_{i,1}$ is the prediction for the target task. We share PLM weights for three types of medical codes to improve efficiency.
Local Model $f_{\theta}$ with Visit Information. To harness the visit-level information, various deep learning architectures have been proposed. In principle, $g_{\phi}$ can be combined with any $f_{\theta}$ to improve performance. In main experiments, we use a hypergraph transformer (HyGT, Xu et al. (2022); Cai et al. (2022)) due to its strong ability to capture high-order relationships between visits and medical codes. It first builds hypergraphs $\mathcal{G}=(\mathcal{V},\mathcal{E})$ by treating medical codes as nodes and patients as hyperedges, then leverages self-attention for aggregating neighborhood information. The details for HyGT are in Appendix C. We obtain the prediction $\widehat{y}_{i,2}$ with $f_{\theta}$ as
$\bm{e}_{i}=\operatorname{HyGT}(\mathcal{G},V_{i}),\widehat{y}_{i,2}=\operatorname{MLP}(\bm{e}_{i}),$ (4)
where $\bm{e}_{i}$ is the representation of patient $i$ after hypergraph transformer.
Co-training. We integrate the two predictors into a co-training framework, with the learning objective:
$\mathcal{L}_{\text{aug}}=$ $\mathbb{E}_{(V_{i},y_{i})\sim\mathcal{P}}~{}\ell(\widehat{y}_{i,1},y_{i})+\lambda\mathcal{D}_{\operatorname{KL}}(\widehat{y}_{i,1},\widetilde{y}),$ (5)
$\mathcal{L}_{\text{loc}}=$ $\mathbb{E}_{(V_{i},y_{i})\sim\mathcal{P}}~{}\ell(\widehat{y}_{i,2},y_{i})+\lambda\mathcal{D}_{\operatorname{KL}}(\widehat{y}_{i,2},\widetilde{y}),$
where $\ell(\cdot)$ is the binary cross-entropy loss, $\widetilde{y}=\beta\widehat{y}_{i,1}+(1-\beta)\widehat{y}_{i,2}$, $\lambda,\beta$ are two hyperparameters. Two losses in Eq. 5 are designed to encourage $f_{\theta}$ and $g_{\phi}$ regularize each other, which can stabilize the learning for two models. During the inference stage, we directly use the $\widetilde{y}_{j}$ as the final prediction for the $j$-th test example $p_{j}$. The overall algorithm of Ram-EHR is illustrated in Algorithm 1.
4 Experiments
4.1 Experiment Setups
Datasets. We conduct experiments on the public MIMIC-III dataset (Johnson et al., 2016) and a private Ih-D1 dataset collected from a large healthcare system in the United States. We perform a 25-label phenotypes prediction task on MIMIC-III, and a cardiovascular disease (CVD) endpoints prediction task for diabetes patients on Ih-D1. We randomly split them into train/validation/test sets by 7:1:2. Please refer to Appendix A for details.
Table 1: Performance on two EHR datasets compared with baselines. The result is averaged over five runs. We use * to indicate statistically significant results ($p<0.05$). For ‘w/o Retrieval’, we directly use LLM to generate summarized knowledge. For ‘w/o LLM Summarization’, we concatenate top-$k$ retrieved documents as summarized knowledge. ‘w/ $g_{\phi}$ only’ means we set $\lambda=0,\beta=1$ (i.e., only use the prediction from $g_{\phi}$ as the final prediction).
Model MIMIC-III Ih-D1
ACC AUROC AUPR F1 ACC AUROC AUPR F1
Transformer (Li et al., 2020) 76.18 80.61 67.12 42.75 78.10 69.49 40.14 58.23
GCT (Choi et al., 2020) 77.20 78.62 64.87 37.57 76.51 68.31 37.55 44.10
HyGT (Cai et al., 2022) 78.07 81.09 68.08 44.93 79.45 70.59 41.04 60.00
MedRetriever (Ye et al., 2021) 77.15 80.14 68.45 39.29 78.95 70.07 42.19 57.96
SeqCare (Xu et al., 2023b) 77.44 80.98 68.53 39.16
GraphCare (Jiang et al., 2024) 80.11 82.26 71.19 44.33 79.09 71.12 43.98 59.00
CORE (van Aken et al., 2021) 79.63 82.05 70.79 43.76 77.11 67.84 40.74 61.12
BEEP (Naik et al., 2022) 79.90 82.67 71.58 44.15 79.29 68.59 41.93 60.95
Ram-EHR 81.59* (1.8%) 84.97* (2.8%) 74.64* (4.3%) 48.19* (7.2%) 80.41* (1.2%) 73.80* (3.8%) 48.40* (10.1%) 63.98* (4.7%)
     w/o Retrieval 80.68 83.29 72.95 44.65 79.83 73.06 47.05 63.25
     w/o LLM Summarization 80.08 82.14 71.35 41.49 77.30 69.71 42.58 61.70
     w/ Augmented Model $g_{\phi}$ Only 81.04 83.80 73.41 46.83 79.70 73.15 47.62 63.33
Evaluation Metrics. Following Choi et al. (2020), we employ Accuracy, AUROC, AUPR, and Macro-F1 as evaluation metrics, where AUROC is the main metric. For accuracy and F1 score, we use a threshold of 0.5 after obtaining predicted results.
Baselines. We consider the following baselines, with a focus on knowledge-enhanced EHR predictive models:
  • Transformer (Li et al., 2020): It leverages the Transformer (Vaswani et al., 2017) architecture to model sequential EHR visits for clinical prediction tasks.
  • GCT (Choi et al., 2020): It employs the Transformer model to learn the EHR’s hidden structure via medical codes. Additionally, it introduces the Graph Convolutional Transformer, integrating graph neural networks to utilize the EHR structure for prediction.
  • HyGT (Cai et al., 2022): It leverages hypergraph transformers that regard patients as hyperedges and medical codes as nodes for EHR predictive tasks.
  • MedRetriever (Ye et al., 2021): It retrieves the most relevant text segments from a local medical corpus using string similarity. Then, it uses query features aggregated with EHR embeddings and disease-specific documents via self-attention for disease prediction tasks.
  • SeqCare (Xu et al., 2023b): It is one of the strongest models for knowledge-graph-based EHR prediction with pruned graph structure learning and debiased training techniques.
  • GraphCare (Jiang et al., 2024): It generates personalized knowledge graphs via prompting LLMs and leverages attention-based graph neural networks for healthcare predictions.
  • CORE (van Aken et al., 2021): It integrates clinical knowledge with specialized outcome pre-training, and uses language models to predict clinical notes for prediction.
  • BEEP (Naik et al., 2022): It augments the language models with the retrieved PubMed articles and fuses them with information from notes to predict clinical outcomes.
(a) Effect of $g_{\phi}$ on MIMIC-III
(b) Effect of $g_{\phi}$ on Ih-D1
(c) Effect of $f_{\theta}$ on MIMIC-III
(d) Effect of $f_{\theta}$ on Ih-D1
Figure 2: Effect of $g_{\phi}$ and $f_{\theta}$ on both datasets.
Implementation Details. In this work, we use Dragon (Lin et al., 2023) as the retriever for both passage $R_{D}(\cdot)$ and queries $R_{Q}(\cdot)$ [^1] https://huggingface.co/facebook/dragon-plus-{query,context}-encoder. We set $k=5$ during the retrieval stage. We choose UMLS-BERT (Michalopoulos et al., 2021) as the augmented model $g_{\phi}$ with 110M parameters for Ram-EHR and relevant baselines as $g_{\phi}$, with a maximum length of 512, and HyGT (Cai et al., 2022) as $f_{\theta}$ in main experiments. We also evaluate the effects of $g_{\phi}$ on PubmedBERT (Tinn et al., 2021), Clin-MobileBERT, and Clin-DistilBERT (Rohanian et al., 2023), as well as the effects of $f_{\theta}$ on GCT (Choi et al., 2020) and Transformer (Li et al., 2020) in Section 4.3. We set the learning rate to 5e-5 for $g_{\phi}$ and 1e-4 for $f_{\theta}$, batch size to 32, and the number of epochs to 5. We select $\beta,\lambda$ based on the performance of the validation set. All the experiments are conducted on a Linux server with one NVIDIA A100 GPU.
4.2 Main Experimental Results
Table 1 exhibits the experiment results of Ram-EHR and baselines. First, we observe Ram-EHR surpasses baselines lacking external knowledge, highlighting the benefits of retrieval augmentation. Second, Ram-EHR outperforms knowledge-enhanced baselines due to the diverse collection of external knowledge as well as the co-training scheme that leverages information from both visit and semantic perspectives. Third, directly using medical notes leads to inferior outcomes due to potential irrelevance, whereas combining medical codes with summarized knowledge as Ram-EHR proves more effective for prediction tasks.
4.3 Additional Studies
4.3.1 Ablation Study
On the bottom of Table 1, we inspect different components in Ram-EHR and observe that removing any of them hurts the performance, which justifies the necessity of our designs. As an illustration, eliminating the retrieval module might result in less informative summarized knowledge. Conversely, excluding the LLM summarization module (i.e., relying solely on retrieved text) could yield lengthy outputs and fail to incorporate all the condensed knowledge about medical codes into the augmented model. Besides, we observe that using the summarized knowledge with $g_{\phi}$ already achieves strong performance, highlighting the benefit of capturing the semantics of medical codes.
(a) Effect of $\mathcal{M}$, MIMIC-III
(b) Effect of $\mathcal{M}$, Ih-D1
Figure 3: Studies on Information Source $\mathcal{M}$.
4.3.2 Effect of $f_{\theta}$ and $g_{\phi}$
We evaluate the effect of the augmented model $g_{\phi}$ and the local model $f_{\theta}$ on both datasets in Figure 2. The experimental results demonstrate that both models contribute to the performance gain. Moreover, it is observed that Ram-EHR is flexible to be applied upon different models, with a comparable performance. Notably, even with a lightweight $f_{\theta}$ (Clin-MobileBERT) having only 25M parameters, Ram-EHR reaches close performance to UMLS-BERT, providing an efficient option for EHR predictive modeling.
4.3.3 Effect of Information Source $\mathcal{M}$
We then evaluate the effectiveness of each knowledge source within $\mathcal{M}$. Figure 3 indicates that incorporating all corpus yields the highest performance, highlighting the value of diverse corpora. Besides, using Drugbank alone contributes minimally, likely due to its limited scope of medication information. Moreover, we observe that leveraging knowledge bases (e.g., MeSH) is more beneficial than literature sources, as they offer broader and more generic information for EHR prediction tasks.
(a) $\beta$ in Eq. 5
(b) $\lambda$ in Eq. 5
Figure 4: Parameter studies of $\beta$ and $\lambda$ on both datasets.
(a) Case Study
(b) Human Study
Figure 5: Case study and human study. The case study compares knowledge summarized by Ram-EHR and directly generated by LLM prompting. Bold denotes disease, medication and procedure concepts. Blue and Red indicate useful and irrelevant knowledge.
4.3.4 Parameter Study
In Figure 4, we conduct parameter studies on both datasets for $\beta$ and $\lambda$ in Eq. 5. Figure 4(a) demonstrates that the model achieves the best performance when $\beta$ is set to 0.2 and 0.4 on MIMIC-III and Ih-D1, respectively, while the gain diminishes at the extremes. This highlights the contribution of combining the predictions from both the augmented model and the local model on the performance gain. In addition, $\lambda$ is set to 1 and 5 on MIMIC-III and Ih-D1, respectively, according to Figure 4(b). The positive values of $\lambda$ indicate that the consistency loss enhances model performance.
4.4 Case Study
Figure 5(a) presents a case study on Ih-D1 to compare knowledge summarized by Ram-EHR and directly generated by LLM prompting. We observe that Ram-EHR provides more relevant information for the downstream task, particularly regarding the CVD outcome in this case, compared to direct LLM prompting. This also aligns with the human study evaluating the quality of 40 randomly sampled knowledge per type of code on a scale of [0,1,2] in Figure 5(b). Specifically, a score of 0 represents irrelevant knowledge, 1 denotes somewhat useful knowledge and 2 indicates highly useful knowledge. The detailed guidelines provided for the annotators are presented in Appendix E.
5 Conclusion
We propose Ram-EHR, which uses dense retrieval with multiple knowledge sources and consistency regularization to enhance EHR prediction tasks. Experiments on two EHR datasets show the efficacy of Ram-EHR over baselines with a gain of 3.4% in AUROC and 7.2% in AUPR. Human studies confirm the usefulness of generated knowledge.
Acknowledgement
We thank the anonymous reviewers and area chairs for valuable feedbacks. This research was partially supported by the internal funds and GPU servers provided by the Computer Science Department of Emory University. JH was supported by NSF grants IIS-1838200 and IIS-2145411.
Limitations
In this work, we propose Ram-EHR to unify external knowledge in text format and adapt it for EHR predictive tasks. Despite its strong performance, we have listed some limitations of Ram-EHR:
Building Multi-source Corpus $\mathcal{M}$.
In this study, we construct a multi-source corpus $\mathcal{M}$ by manually selecting five relevant sources within the clinical domain. In real-world scenarios, the grounding corpora usually require customization according to query domains and user needs. Therefore, effectively selecting grounding corpora and efficiently evaluating their relative contributions remains an unresolved issue. Furthermore, retrieved evidence may contain noise that could potentially degrade model performance, highlighting the importance of developing fine-grained filtering or re-ranking modules as a crucial area for future research.
Other Types of Models and Tasks.
Although we have tested three $f_{\theta}$ and four $g_{\phi}$ combinations on two clinical prediction tasks, there are numerous other clinical tasks and prediction models that exist. Due to computational constraints, we are unable to explore all possible model combinations. Extending Ram-EHR to encompass additional models and tasks is a crucial avenue for future research.
Efficiency.
The integration of the augmented model $g_{\phi}$ can result in additional time complexity. In our main experiment setups (using UMLS-BERT), co-training usually takes 1.5×1.5$\times$1.5 × to 2×2$\times$2 × more times than using the local model alone. One potential solution is to use a lightweight model (e.g., Clin-MobileBERT) to improve efficiency.
Ethical Considerations
One potential ethical consideration concerns the use of credential data (MIMIC-III and Ih-D1) with GPT-based online services. We have signed and strictly adhered to the PhysioNet Credentialed Data Use Agreement [^2] https://physionet.org/about/licenses/physionet-credentialed-health-data-license-150/ for the legal usage of the MIMIC-III dataset. To prevent sensitive information from being shared with third parties through APIs, we carefully follow the guidelines [^3] https://physionet.org/news/post/gpt-responsible-use for the responsible use of MIMIC data in online services. Specifically, we have requested to opt out of human review of the data by filling out the Azure OpenAI Additional Use Case Form [^4] https://aka.ms/oai/additionalusecase in order to utilize the Azure Open AI service while ensuring that Microsoft does not have access to the patient data. The utilization of LLMs in our framework is strictly to build medical concept-specific KGs. In addition, the building of medical concept-specific KGs does not involve direct interaction with any individual patient information. We iterate through all concepts in the medical coding system (e.g., CCS and ICD) to generate their respective KGs using LLMs, and these KGs are stored locally.
Appendix A Task Information
We present the detailed statistics of MIMIC-III and Ih-D1 in Table 3.
MIMIC-III.
The MIMIC-III dataset (Johnson et al., 2016) is a large, freely available database that contains de-identified health-related data from over 4,000 patients who stayed in critical care units at the Beth Israel Deaconess Medical Center between 2001 and 2012. We conduct the phenotyping prediction task proposed by Harutyunyan et al. (2019). It aims to predict whether the 25 pre-defined acute care conditions (see Table 2) are present in a patient’s next visit, based on the information from their current visit. The problem is formulated as a 25-label binary classification, considering that multiple phenotypes may exist in a single visit. For data preprocessing, we focus on patients with multiple hospital visits, identified based on their admission information. We extract pairs of consecutive visits for each patient. For each pair, we extract diseases, medications, and procedures from the health records in the former visit as input, and identify the phenotypes in the latter visit as labels, using Clinical Classifications Software (CCS) from the Healthcare Cost and Utilization Project (HCUP) [^5] https://hcup-us.ahrq.gov/toolssoftware/ccs/AppendixASingleDX.txt.
Table 2: The 25 pre-defined phenotypes in MIMIC-III.
Phenotype Type
Acute and unspecifed renal failure acute
Acute cerebrovascular disease acute
Acute myocardial infarction acute
Cardiac dysrhythmias mixed
Chronic kidney disease chronic
Chronic obstructive pulmonary disease chronic
Complications of surgical/medical care acute
Conduction disorders mixed
Congestive heart failure; nonhypertensive mixed
Coronary atherosclerosis and related chronic
Diabetes mellitus with complications mixed
Diabetes mellitus without complication chronic
Disorders of lipid metabolism chronic
Essential hypertension chronic
Fluid and electrolyte disorders acute
Gastrointestinal hemorrhage acute
Hypertension with complications chronic
Other liver diseases mixed
Other lower respiratory disease acute
Other upper respiratory disease acute
Pleurisy; pneumothorax; pulmonary collapse acute
Pneumonia acute
Respiratory failure; insufficiency; arrest acute
Septicemia (except in labor) acute
Shock acute
Ih-D1.
For the Ih-D1 dataset, we conduct a CVD outcome prediction task, which predicts whether patients with type 2 diabetes will experience CVD complications within 1 year after their initial diagnosis, including coronary heart disease (CHD), congestive heart failure (CHF), myocardial infarction (MI), or stroke. Diseases are identified by their ICD-9 or ICD-10 clinical codes. The usage of data has been approved by the Institutional Review Board (IRB).
Table 3: The statistics of MIMIC-III and Ih-D1.
Stats MIMIC-III Ih-D1
# of diagnosis 846 7915
# of medication 4525 489
# of procedure 2032 4321
# of health records 12353 36611
Appendix B Knowledge Sources
B.1 Descriptions
  • PubMed [^6] https://pubmed.ncbi.nlm.nih.gov/: PubMed is a free search engine accessing primarily the MEDLINE database of references and abstracts on life sciences and biomedical topics. It provides users with access to millions of scientific documents, including research papers, reviews, and other scholarly articles. We use the Entrez package to extract the PubMed articles [^7] https://biopython.org/docs/1.75/api/Bio.Entrez.html, resulting in 230k documents.
  • DrugBank [^8] https://go.drugbank.com/releases/latest (Wishart et al., 2008): DrugBank is a comprehensive and freely accessible online database containing information on drugs and drug targets. It integrates detailed drug data (chemical, pharmacological, and pharmaceutical) with comprehensive information on drug targets (sequence, structure, and pathway). We use the data from the original database, which contains 355k documents.
  • Medical Subject Headings (MeSH) [^9] https://www.ncbi.nlm.nih.gov/mesh/: Medical Subject Headings (MeSH) is a comprehensive controlled collection for indexing journal articles and books in the life sciences. It organizes information on biomedical and health-related topics into a hierarchical structure. The corpus contains 32.5k documents covering various medical concepts.
  • Wikipedia [^10] https://www.wikipedia.org/ (Vrandečić and Krötzsch, 2014): Wikipedia is a free, web-based, collaborative, multilingual encyclopedia project that is supported by the non-profit Wikimedia Foundation. We extract web pages that contain medical-related information by using the medical codes list (e.g., ICD10 and ATC), resulting in 150k documents.
  • KG [^11] https://github.com/mims-harvard/PrimeKG (Chandak et al., 2023): We use PrimeKG in our experiments. It offers a comprehensive overview of diseases, medications, side effects, and proteins by merging 20 biomedical sources to detail 17,080 diseases across ten biological levels. For this study, we select knowledge triplets that contain medical codes within three types (disease, medication, procedure) used in this work, resulting in 707k triplets. We use the template in Appendix B.2 to transform these triplets into sentences.
B.2 Translating Format
We list the template to transform knowledge triplets into sentences in KG as follows:
candidate_relation = ["disease_phenotype_positive", "disease_protein", "disease_disease", "drug_effect", "drug_protein"]
relations = {
"phenotype present": "[ent1] has the phenotype [ent2]",
"carrier": "[ent1] interacts with the carrier [ent2]",
"enzyme": "[ent1] interacts with the enzyme [ent2]",
"target": "The target of [ent1] is [ent2]",
"transporter": "[ent2] transports [ent1]",
"associated with": "[ent2] is associated with [ent1]",
"parent-child": "[ent2] is a subclass of [ent1]",
"side effect": "[ent1] has the side effect of [ent2]"
Appendix C Details for Hypergraph Transformer
First of all, we construct a hypergraph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ based on EHR data, where each patient visit is represented as a hyperedge connecting to all medical codes associated with the visit as nodes. Then we utilize HyGT (Cai et al., 2022) to jointly learn the node and hyperedge embeddings. Specifically, The hyperedge embeddings aggregate information from nodes within each hyperedge, while the node embeddings aggregate information from hyperedges connecting the nodes. In the $l$-th neural network layer, the node and hyperedge embeddings are updated as
$\ \bm{X}_{v}^{(l)}=f_{\mathcal{E}\rightarrow\mathcal{V}}\left(\mathcal{E}_{v,\bm{E}^{(l-1)}}\right),$ (6)
$\bm{E}_{e}^{(l)}=f_{\mathcal{V}\rightarrow\mathcal{E}}\left(\mathcal{V}_{e,\bm{X}^{(l-1)}}\right),$ (7)
where $\bm{X}_{v}^{(l)}$ and $\bm{E}_{e}^{(l)}$ represent the embeddings of node $v$ and hyperedge $e$ in the $l$-th layer ($1\leq l\leq L$), respectively. $\mathcal{E}_{v,\bm{E}}$ denotes the hidden representations of hyperedges that connect the node $v$, while $\mathcal{V}_{e,\bm{X}}$ is the hidden representations of nodes that are contained in the hyperedge $e$. The two message-passing functions $f_{\mathcal{V}\rightarrow\mathcal{E}}(\cdot)$ and $f_{\mathcal{E}\rightarrow\mathcal{V}}(\cdot)$ utilize multi-head self-attention (Vaswani et al., 2017) to identify significant neighbors during propagation as
$f_{\mathcal{V}\rightarrow\mathcal{E}}(\bm{S})=f_{\mathcal{E}\rightarrow\mathcal{V}}(\bm{S})=\operatorname{Self-Att}(\bm{S}),$
[…]
$\text{SA}_{i}(\bm{S})=\text{softmax}\left(\frac{\bm{W}_{i}^{Q}(\bm{SW}_{i}^{K})^{\top}}{\sqrt{\lfloor d/h\rfloor}}\right)\bm{SW}_{i}^{V}.$
[…]
Appendix D Details for Prompt Design
We present the detailed design of the prompt template as follows:
 Prompt for LLM Summarization
 Suppose you are a physician working on a health
     -related outcome prediction task and need
     to get relevant information for the given
     <task>. Here is relevant information:
 <medical code type> Name: <medical code name>
 Retrieve Passage #1: <retrieved document 1>
 Retrieve Passage #2: <retrieved document 2>
 Based on the above information, Could you
     generate 1 sentence of around 10-20 words
     to summarize the knowledge for the
     <medical code type> that is useful for
     the <task>?     <task> is the brief description of the downstream task. <medical code type> is either “disease”, “medication” or “procedure”, depending on the input.
Appendix E Human Study Guidelines
For the human studies in Section 4.3, we provide the following guidelines for the annotators to evaluate the quality of the generated knowledge.
The goal of this evaluation is to assess the helpfulness of generated knowledge explaining or relating to specific medical codes in the context of target prediction tasks. Helpfulness is defined by the relevance, accuracy, and utility of the information in facilitating understanding or decision-making related to medical coding and its implications for predictive tasks.
Please rate the following generated knowledge with score 0, 1 or 2.
> 0: Irrelevant
Definition: The knowledge does not provide any relevant information related to the medical code in question. It might be factually accurate but completely off-topic or not applicable to the context of target prediction tasks.
> 1: Partially Relevant and Useful
Definition: The knowledge provides some relevant information but either lacks completeness, specificity, or direct applicability to target prediction tasks. It might include general facts or insights that are related to the medical code but does not fully support decision-making or understanding in a predictive context.
> 2: Very Useful
Definition: The knowledge directly addresses the medical code with accurate, relevant, and comprehensive information that is highly applicable to target prediction tasks. It should provide detailed understanding, or specific examples that facilitate decision-making, understanding, or application in predictive modeling.
Appendix F Cost Information
Utilizing GPT-3.5-turbo as our base LLM model for generating summarized knowledge, we observe an average retrieval augmentation cost of \$0.0025 per medical code in MIMIC-III and \$0.0032 in Ih-D1. Consequently, Ram-EHR does not result in excessive monetary expenses.